# Joint Multimodal Transformer for Emotion Recognition in the Wild

### v2

@@ 5.1 Comparison with the State-of-the-Art
Table 1 presents the performance of our fusion model on the Biovid Heat Pain Database. We performed 5-fold cross-validation to pick up the best average fusion model. It can be seen from the table that the proposed model can achieve state-of-the-art performance with multimodal input while using 5-fold cross-validation. The physiological modality is stronger in the Biovid database. Many studies have validated the models on physiological modality. Our empirical results show that the EDA-only accuracy is 77.2%, whereas the visual-only accuracy is 72.9%. The proposed model can improve over unimodal performance and achieves state-of-the-art performance on the Biovid dataset. We also compare it with standard fusion techniques like feature concatenation. For a fair comparison, we keep all the parameters the same. The proposed model improves 6% over simple feature concatenation and 1.3% over a vanilla multimodal transformer i.e. without a joint representation. Figure 4 shows the visualization of the attention weights generated by the joint transformer model. It can be seen that the model gives more weightage to the physiological modality.
@@
Table 3 shows the valence and arousal CCC values on the Affwild2 official validation set and custom-defined folds to increase generalizability. On the official validation set, the proposed method achieves a 0.666 average with 0.717 valence and 0.614 arousal.
@@
Table 4 compares the proposed method with the state-of-the-art on Affwild2 development set. Our method achieves a 0.722 average with 0.76 valence and 0.68 arousal.
Table 4: CCC performance of the proposed and state-of-the-art methods for A-V fusion on the Affwild2 development set.
@@
Joint Multimodal Transformer (ours) 0.760 0.684 0.722
Situ-RUCAIM3 [37] 0.655 0.709 0.682
Joint Cross- Attention [46] 0.670 0.590 0.630
FlyingPigs [62] 0.559 0.671 0.615
PRL [41] 0.390 0.540 0.465
AU-NO [14] 0.361 0.551 0.456
HSE-NN [49] 0.408 0.477 0.443
Table 5 compares the proposed method with the state-of-the-art on the Affwild2 test set. The proposed method can achieve 0.42 valence and 0.46 arousal. The average of valence and arousal is 0.443. The proposed method significantly improves over the baseline, which is 0.17 average. It is essential to mention here that the other methods that achieve higher performance are due to extensive pertaining, the use of additional modalities like text, and the use of more robust backbones. For a fair comparison, we use a similar setting to Praveen et al. [46], which includes similar pretraining and identical backbones for audio and visual modalities. They used joint cross-attention to fuse the two modalities and achieve a 0.369 average. On the other hand, the proposed model uses a joint multimodal transformer-based fusion method and can achieve a 0.443 average. The proposed method improves 7% over the joint cross-attention-based method.
Table 5: CCC performance of the proposed method and state-of-the-art methods for A-V fusion on the Affwild2 test set.
Method Valence Arousal Mean
@@
FlyingPigs [62] 0.520 0.602 0.561
PRL [41] 0.450 0.445 0.448
Joint Multimodal Transformer (ours) 0.420 0.467 0.443
HSE-NN [49] 0.417 0.454 0.436
AU-NO [14] 0.418 0.407 0.413
Joint Cross- Attention [46] 0.374 0.363 0.369
Baseline 0.180 0.170 0.175
@@ 5.2 Ablations
To see the effectiveness of the joint representation of the proposed method, we performed an ablation study. Table 6 shows the results of the proposed method with and without the joint representation. On the Biovid dataset, the joint multimodal transformer improves by 1.3% over the vanilla multimodal transformer.
@@
Table 7 shows that the proposed model improves accuracy over the vanilla multimodal transformer by 1.8% with a R2D1 and ResNet18 backbone on the Affwild2 dataset. With an I3D and a ResNet18 backbone, results are improved by 1.7%, as shown in Table 8. Notice that the results on the test set are different from Table 5 because we used a different train/validation split.
Fusion Model Valence Arousal Mean
R2D1 model only (visual) 0.194 0.310 0.252
ResNet18 model only (audio) 0.273 0.246 0.260
Concat + FC layers 0.320 0.327 0.323
Vanilla Transformer 0.376 0.334 0.355
Joint Multimodal Transformer (ours) 0.366 0.379 0.373
Table 8: CCC performance of the backbones alone, and combined using baselines and the proposed A-V fusion method on the Affwild2 test set. Experiments are performed on the default training-validation split. I3D (visual) and ResNet18 (audio) backbones are used for all cases.
Fusion Model Valence Arousal Mean
I3D model only (visual) 0.336 0.422 0.379
ResNet18 model only (audio) 0.273 0.246 0.260
Concat + FC layers 0.387 0.453 0.420
Vanilla Transformer 0.432 0.410 0.421
Joint Multimodal Transformer (ours) 0.425 0.450 0.438
Joint Cross Attention [46] 0.374 0.363 0.369
@@ 6 Conclusion
Multimodal emotion recognition systems outperform their unimodal counterparts, especially in the wild environment. The missing and noisy modality is a prevalent issue with in-the-wild emotion recognition systems. Many attention-based methods have been proposed in the literature to overcome this problem. These methods aim to weigh the modalities dynamically. This paper introduces a joint multimodal transformer for dimensional emotional recognition in the wild. This transformer-based architecture introduces a joint feature representation in order to add more redundancy and complementary between audio and visual data. The proposed model outperforms the state of the art on the Biovid dataset and improves 6% on the Affwild2 dataset. In the future, we aim to introduce more modalities and sophisticated backbones for effective feature extraction.

### v3

@@ 5.1 Comparison with the State-of-the-Art
This section compares the proposed method with the baseline and state-of-the-art on the Affwild2 and Biovid datasets.
Table 2: A comparison of the proposed JMT and state-of-the-art fusion models for pain estimation on the Biovid Heat Pain Database. The highest score is indicated in bold.
Method Modality XV Scheme Accuracy (%)
Werner et al. [60] ICPR 2014 EDA, ECG, EMG + Video 5-FOLD 80.6
Werner et al. [58] IEEE TAC 2016 Video LOSO 72.4
Kachele et al. [12] IEEE IJSTSP 2016 EDA, ECG, EMG LOSO 82.73
Lopez et al. [33] ACIIW 2017 EDA, ECG 10-FOLD 82.75
Lopez et al. [34] EMBC 2018 EDA LOSO 74.21
Thiam et al. [51] Sensors 2019 EDA LOSO 84.57
Wang et al. [56] EMBC 2020 EDA, ECG, EMG LOSO 83.3
Pouromran et al. [45] PLoS ONE 2021 EDA LOSO 83.3
Thiam et al. [52] Frontiers 2021 EDA, ECG, EMG LOSO 84.25
Phan et al. [44] IEEE Access 2023 EDA, ECG, EMG LOSO 84.8
Audio backbone: 1D CNN only EDA 5-FOLD 77.2
Visual backbone: R3D model only Video 5-FOLD 72.9
Fusion: feature concatenation EDA + Video 5-FOLD 83.5
Fusion: vanilla transformer EDA + Video 5-FOLD 87.8
Fusion: JMT (ours) EDA + Video 5-FOLD 89.1
Table 2 presents the performance of our fusion model on the Biovid Heat Pain Database. We performed 5-fold cross-validation to pick up the best average fusion model. It can be seen from the table that the proposed model can achieve state-of-the-art performance with multimodal input while using 5-fold cross-validation. We used 5-fold cross-validation instead of LOSO validation method due to the computational cost of processing video. The physiological modality is stronger in the Biovid database. Many studies have validated the models on physiological modality. Our empirical results show that the EDA-only accuracy is 77.2%, whereas the visual-only accuracy is 72.9%. The proposed model can improve over unimodal performance and achieves state-of-the-art performance on the Biovid dataset. We also compare it with standard fusion techniques like feature concatenation. For a fair comparison, we keep all the parameters the same. The proposed model improves 6% over simple feature concatenation and 1.3% over a vanilla multimodal transformer i.e. without a joint representation.
@@
Figure 4 shows the visualization of the attention weights generated by the joint transformer model. It can be seen that the model gives more weightage to the physiological modality.
Table 3 shows the valence and arousal CCC values on the Affwild2 official validation set and custom-defined folds to increase generalizability. On the official validation set, the proposed method achieves a 0.666 average with 0.717 valence and 0.614 arousal. In addition, we performed ensembling by considering a dual modeling. We observed that different training configurations lead to models that can be best at once case: ’Valence’, or ’Arousal’. Therefore, we used a dual model for prediction by taking the best at each category. This lead to performance increase.
@@
Ensembling 0.769 0.692 0.731
Table 4 compares the proposed JMT fusion method against state-of-the-art fusion methods on the Affwild2 test set. The proposed method can achieve 0.472 valence and 0.443 arousal. The average of valence and arousal is 0.458. The proposed method significantly improves over the baseline (provided by the challenge organizers). It is essential to mention here that the other methods that achieve higher performance are due to extensive pertaining, the use of additional modalities like text, and the use of more robust backbones. For a fair comparison, we use a similar setting to Joint Cross Attention [46], which includes similar pretraining and identical backbones for audio and visual modalities. They used joint cross-attention to fuse the two modalities and achieve a 0.369 average. On the other hand, the proposed model uses JMT fusion and can achieve a 0.443 average. The proposed method improves 7% over the joint cross-attention-based method.
Table 4: CCC performance of the proposed JMT and state-of-the-art methods for A-V fusion on the Affwild2 test set.
@@
Baseline 0.180 0.170 0.175
Joint Cross Attention [46] 0.374 0.363 0.369
AU-NO [14] 0.418 0.407 0.413
HSE-NN [49] 0.417 0.454 0.436
PRL [41] 0.450 0.445 0.448
Joint Multimodal Transformer (ours) 0.472 0.443 0.458
FlyingPigs [62] 0.520 0.602 0.561
@@ 5.2 Ablations
Table 5 shows that the proposed model improves performance over the vanilla multimodal transformer by 1.8% with an R2D1 and ResNet18 backbone on the Affwild2 dataset. With an I3D and a ResNet18 backbone, results are improved by 1.7%, as shown in Table 6. Notice that the results on the test set are different from Table 4 because we used a different train/validation split. Additionally, results show that feature-level fusion of 2 vision backbones (I3D and R2D1) improves the performance compared to using a single vision backbone.
Fusion Model Valence Arousal Mean
R2D1 model only (visual) 0.194 0.310 0.252
ResNet18 model only (audio) 0.273 0.246 0.260
Concat + FC layers 0.320 0.327 0.323
Vanilla Transformer 0.376 0.334 0.355
JMT (ours) 0.366 0.379 0.373
Table 6: CCC performance of visual and audio backbones alone, and their A-V fusion using the proposed and baseline methods on the Affwild2 test set. Experiments are performed on the default training-validation split. I3D (visual) and ResNet18 (audio) backbones are used for all cases. FC: fully connected layers for feature-level fusion of backbones from the same modality. TR: transformer for fusing features from backbones of the same modality.
Fusion Model Valence Arousal Mean
Joint Cross Attention [46] 0.374 0.363 0.369
I3D model only (visual) 0.336 0.422 0.379
ResNet18 model only (audio) 0.273 0.246 0.260
Concatenation + FC 0.387 0.453 0.420
Vanilla Transformer - I3D (visual) - ResNet18 (audio) 0.432 0.410 0.421
JMT (ours) - I3D (visual) - ResNet18 (audio) 0.425 0.450 0.438
JMT (ours) - I3D+R2D1 (visual, FC) - ResNet18 (audio) 0.472 0.443 0.458
JMT (ours) - I3D+R2D1 (visual, TR) - ResNet18 (audio) 0.458 0.445 0.452
Table 7 shows the results of the proposed method with and without the joint representation. On the Biovid dataset, the joint multimodal transformer improves by 1.3% over the vanilla multimodal transformer.
@@ 6 Conclusion
Multimodal emotion recognition systems outperform their unimodal counterparts, especially in the wild environment. The missing and noisy modality is a prevalent issue with in-the-wild emotion recognition systems. Many attention-based methods have been proposed in the literature to overcome this problem. These methods aim to weigh the modalities dynamically. This paper introduces a joint multimodal transformer for emotional recognition. This transformer-based architecture introduces a joint feature representation to add more redundancy and complementary between audio and visual data. The two modalities are first encoded using separate backbones to extract intra-modal spatiotemporal dependencies. The feature vectors of the two modalities are joint and the joint feature vector is also fed into the Joint Multimodal Transformer module. This joint representation provides more fine-grained information about the inter-modal association between the two modalities. The proposed model outperforms state-of-the-art methods on the Biovid dataset and improves over the vanilla multimodal transformer by 6% on the Affwild2 dataset. Our future work includes introducing more modalities and sophisticated backbones for effective feature extraction.
Acknowledgement
This work was supported by the Fonds de recherche du Québec – Santé (FRQS), the Natural Sciences and Engineering Research Council of Canada (NSERC), and Canada Foundation for Innovation (CFI).